Parth (parthkohli):

Challenge! If \(a, b, c \in \mathbb{R}\) and \(c(a+b+c)<0\), then the quadratic equation \(ax^2 + bx + c = 0\) definitely has: A. A negative root. B. Two complex roots. C. Two real roots. D. Two positive real roots.

hero (hero):

I predict that it is D

Parth (parthkohli):

Please show the solution along with your answer.

hero (hero):

I assume a and b are negative and c remains positive. And (a + b) < c This creates the scenario where c(a + b + c) < 0 Next I utilize b^2 - 4ac Since a and b is negative (-b)^2 - 4(-a)(c) = b^2 + 4ac which means the result will be positive As long as a + b < c then the result will always yield positive discriminant.

OpenStudy (anonymous):

a = b = -2; c = 1; We have the following roots \[ \left\{\left\{x\to \frac{1}{2} \left(-1-\sqrt{3}\right)\right\},\left\{x\to \frac{1}{2} \left(\sqrt{3}-1\right)\right\}\right\} \]

Parth (parthkohli):

Positive discriminant \(\ne\) positive roots.

OpenStudy (kc_kennylau):

What if c is negative and a+b>|c| ?

hero (hero):

Correction: The answer is C. If you have a positive discriminant, then there will always be two real roots.

OpenStudy (anonymous):

simply f(0). f(1) < 0 use location of roots to predict its nature.... and one of the root will be negative

Parth (parthkohli):

Bansal Classes. :D By the way, what if the graph is downwards? Still negative?

OpenStudy (anonymous):

yep...bansal still the same...

Parth (parthkohli):



Parth (parthkohli):

I've almost given off the solution. So I'll better do the whole thing now. [SPOILER ALERT] @divu.mkr followed the right path. \[c(a + b + c) < 0\]\[\Rightarrow f(0) \times f(1) < 0\]So \(f(0)\) and \(f(1)\) have the opposite sign. So we have a root lying between \(0\) and \(1\). For obvious reasons, the other root is real too. And so we're left with C as @Hero said.

Parth (parthkohli):

Just noticed that the question has a gap.

ganeshie8 (ganeshie8):

Very cool(unexpected) way to use IVT ! xD

Parth (parthkohli):

Haha, yeah. I was stunned at the solution. By the way, err, this expression can be a perfect square too.